OpenStudy (anonymous):

Above are two different models of a microchip. If the model of the microchip on the left measures 6 cm tall, how tall is the model of the microchip on the right? 16 cm 10 cm 6 cm 12 cm

OpenStudy (anonymous):

model one is 1cm=1/5mm Model two is 1cm= 1/10mm

OpenStudy (anonymous):

please help @toodles and @bananas1234

OpenStudy (anonymous):

help @byronichero

OpenStudy (anonymous):

12 i think

OpenStudy (anonymous):

thxs ill try @byronichero

OpenStudy (anonymous):

it was right. medal u get